OpenStudy (anonymous):

If f(x) = 3x + 1, then f(a + h) - f(a) = A.h B.3h C.3h + 2

OpenStudy (callisto):

f(a + h) - f(a) = 3(a+h) + 1 - (3a+1) = 3h